# The Open Journal of Astrophysics

### v1

@@ 6.2 Correlation Function Objective
where $\sigma^{2}_{\omega}(r_{k})$ is the empirical variance in bin $k$ and $\epsilon=10^{-6}$ provides numerical stability to prevent division by zero in bins with very low variance. The weights are normalized such that $\sum_{k}W_{k}=1$. This inverse-variance weighting approximates the covariance due to galaxy shape noise, which is in general dominant over the sample variance at the scales being considered (Van Alfen et al., 2024). While the jackknife covariances that account for both sample variance and galaxy shape noise from the TNG300 data itself are available following the procedure of Van Alfen et al. (2024), we instead use a diagonal variance estimate obtained by averaging over many diffHOD-IA orientation realizations (for fixed galaxy positions), which serves as a proxy for the shape noise and is considerably more computationally efficient. More detail on optimization hyperparameters is given in Appendix C.
@@
where $\mathbf{C}$ is the covariance matrix of $\omega(r)$ estimated from 50 independent orientation samples at the fiducial parameters. We additionally correct for the finite number of realizations by incorporating the Hartlap factor (Hartlap et al., 2006) when inverting the covariance matrix. The scatter in recovered parameters across these 50 catalogs provides an empirical estimate of the covariance, yielding uncertainties $\sigma(\mu_{\text{cen}})=0.008$ and $\sigma(\mu_{\text{sat}})=0.026$, with correlation coefficient $\rho=-0.68$. This anti-correlation aligns with theoretical expectations and was also seen experimentally in Berman et al. (2025). The tighter clustering of converged values reflects the fact that all optimization runs fit the same catalog to the same target, whereas the uncertainty contours capture the full variance including HOD stochasticity.
@@ 6.3 Hamiltonian Monte Carlo
Figure 6: $\mu_{\text{cen}}$ and $\mu_{\text{sat}}$ posteriors for the fiducial tng300 catalog derived using diffHOD-IA with HMC (blue), IAEmu with HMC (orange), and halotools-IA with MCMC (green). diffHOD-IA is in excellent agreement with halotools-IA, while exhibiting substantially faster inference.
Setup. To complement our previous correlation function matching gradient descent experiment, we now turn to HMC to illustrate Monte Carlo inference capabilities with diffHOD-IA. We use the same $\hat{\omega}(r)$ mock observation from the previous section, this time using the jackknife covariance estimate from tng300. For this analysis, we use a fixed HOD and only consider $\mu_{\text{sat}}$ and $\mu_{\text{cen}}$ as free parameters.
@@
Results. Posteriors for diffHOD-IA, halotools-IA, and IAEmu for this experiment are shown in Figure 6. There is in general excellent overlap between the three posteriors, illustrating that diffHOD-IA offers similar inference capabilities as halotools-IA with much faster convergence with HMC. A slight bias is seen in the case of $\mu_{\text{sat}}$, likely propagated from discrepancies at small scales due to satellite occupation. Nonetheless, diffHOD-IA is within $0.4\sigma$ agreement with halotools-IA. Exact posterior values are given in Table 2.
@@ 7 Summary & Discussion
In this work, we have developed diffHOD-IA, a differentiable HOD framework that includes galaxy IA modeling. Our HOD implementation closely follows that of Horowitz et al. (2022), which is extended to the IA modeling of Van Alfen et al. (2024) via inverse Dimroth–Watson CDF sampling. We additionally draw inspiration from Hearin et al. (2022) to extend the differentiability of diffHOD-IA to include IA correlation functions. The diffHOD-IA code is publicly available at .
@@
Results. We benchmarked the accuracy of the sampling procedure and gradients of diffHOD-IA, comparing autodiff computed gradients with analytic and finite-differences, and comparing the differentiable Dimroth Watson samples with samples from the true distribution. diffHOD-IA agreed with halotools-IA across all tests: including galaxy number counts, and the $\xi(r)$, $\omega(r)$, and $\eta(r)$ statistics for the fiducial galaxy catalog.
@@
The differentiability also allows inserting diffHOD-IA into larger differentiable inference pipelines. For example, diffHOD-IA can be integrated into simulation pipelines that incorporate a differentiable particle mesh solver such as jaxPM (Li et al., 2022), along with a differentiable halo finder like jFoF (Horowitz et al., 2022). This would enable end-to-end gradient flow from cosmological, HOD, and IA parameters to the galaxy field.
The differentiable correlation function framework could also be extended to 2D-projected statistics, which are more directly comparable to observational weak lensing measurements (Van Alfen et al., 2025a). Of equal interest is the extension to higher order statistics and field-level inference methods which would impose tighter constraints on parameters. A joint inference over both HOD and IA parameters represents a natural application, using the differentiable estimators for $\xi(r)$ and $\omega(r)$. Such joint analyses over the full 7-dimensional parameter space would be more computationally demanding and are reserved for future work. These developments provide a foundation for efficient, gradient-based joint inference of HOD and IA parameters in current and future weak lensing surveys.
@@ Appendix B Differentiable Galaxy Clustering Statistics
Figure 8: Validation of differentiable correlation function estimators against halotools reference implementations for the same galaxy catalog. Left: Galaxy position-position correlation function $\xi(r)$. The black circles show the (non-differentiable) measurement from halotools.mock_observables.tpcf, while red squares show the differentiable estimator using HOD-derived occupation probability weights. Center: Galaxy position-orientation correlation function $\omega(r)$, comparing halotools.mock_observables.ed_3d (black) with our differentiable estimator (red). Right: Galaxy orientation-orientation correlation function $\eta(r)$, comparing halotools.mock_observables.ee_3d (black) with our differentiable estimator (red). Both $\omega(r)$ and $\eta(r)$ show excellent agreement between implementations, with $\eta(r)$ exhibiting larger statistical fluctuations due to galaxy shape noise. The differentiable estimators enable gradient-based inference: $\xi(r)$ gradients flow through galaxy occupation weights from HOD parameters, while $\omega(r)$ and $\eta(r)$ gradients flow through orientation vectors from IA parameters.

### v2

@@ 6.2 Correlation Function Objective
where $\sigma^{2}_{\omega}(r_{k})$ is the empirical variance in bin $k$ and $\epsilon=10^{-6}$ provides numerical stability to prevent division by zero in bins with very low variance. The weights are normalized such that $\sum_{k}W_{k}=1$. These inverse-variance weights serve only to improve convergence by preventing high-variance bins from dominating the gradient updates, and approximates the covariance due to galaxy shape noise, which is in general dominant over the sample variance at the scales being considered (Van Alfen et al., 2024). While the jackknife covariances that account for both sample variance and galaxy shape noise from the TNG300 data itself are available following the procedure of Van Alfen et al. (2024), we instead use a diagonal variance estimate obtained by averaging over many diffHOD-IA orientation realizations (for fixed galaxy positions), which serves as a proxy for the shape noise and is considerably more computationally efficient. More detail on optimization hyperparameters is given in Appendix C.
@@
where $\mathbf{C}$ is the covariance matrix of $\omega(r)$ estimated from 50 independent orientation samples at the fiducial parameters. We additionally correct for the finite number of realizations by incorporating the Hartlap factor (Hartlap et al., 2006) when inverting the covariance matrix. We emphasize that this empirical covariance is distinct and more accurate than the diagonal variance estimate used in the optimization.
The scatter in recovered parameters across these 50 catalogs provides an empirical estimate of the covariance, yielding uncertainties $\sigma(\mu_{\text{cen}})=0.008$ and $\sigma(\mu_{\text{sat}})=0.026$, with correlation coefficient $\rho=-0.68$. This anti-correlation aligns with theoretical expectations and was also seen experimentally in Berman et al. (2025). The tighter clustering of converged values reflects the fact that all optimization runs fit the same catalog to the same target, whereas the uncertainty contours capture the full variance including HOD stochasticity.
@@ 6.3 Hamiltonian Monte Carlo
Figure 6: $\mu_{\text{cen}}$ and $\mu_{\text{sat}}$ posteriors for the fiducial tng300 catalog derived using diffHOD-IA with HMC (blue), IAEmu with HMC (orange), and halotools-IA with MCMC (green). diffHOD-IA is in excellent agreement with halotools-IA, while exhibiting substantially faster inference. Exact posterior values are given in Table 2.
Setup. To complement our previous correlation function matching gradient descent experiment, we now turn to HMC to illustrate Monte Carlo inference capabilities with diffHOD-IA. We use the same $\hat{\omega}(r)$ mock observation from Section 6.2; however, where a diagonal variance estimate sufficed for gradient-based optimization, here we adopt the jackknife covariance from Pandya et al. (2025) specifically to enable a direct comparison of posteriors between diffHOD-IA, halotools-IA, and IAEmu, all of which used identical covariance estimates in that work. For this analysis, we use a fixed HOD and only consider $\mu_{\text{sat}}$ and $\mu_{\text{cen}}$ as free parameters.
@@
Results. Posteriors for diffHOD-IA, halotools-IA, and IAEmu for this experiment are shown in Figure 6. There is in general excellent overlap between the three posteriors, illustrating that diffHOD-IA offers similar inference capabilities as halotools-IA with much faster convergence with HMC. A slight bias is seen in the case of $\mu_{\text{sat}}$ for both diffHOD-IA and IAEmu compared to halotools-IA. This can potentially be attributed to different seeds in the target galaxy catalog. Nonetheless, diffHOD-IA is within $0.4\sigma$ agreement with halotools-IA, which is consistent with the expected variation for catalogs generated with different random seeds. Exact posterior values are given in Table 2.
@@ 7 Summary & Discussion
In this work, we have developed diffHOD-IA, a differentiable HOD framework that includes galaxy IA modeling. Our HOD implementation closely follows that of Horowitz et al. (2022), which is extended to the IA modeling of Van Alfen et al. (2024) via inverse Dimroth–Watson CDF sampling. We additionally draw inspiration from Hearin et al. (2022) to extend the differentiability of diffHOD-IA to include IA correlation functions. The diffHOD-IA code is publicly available at  [^1] https://github.com/snehjp2/diffHOD-IA.
@@
Results. We benchmarked the accuracy of the sampling procedure and gradients of diffHOD-IA, comparing autodiff computed gradients with analytic and finite-differences, and comparing the differentiable Dimroth–Watson samples with samples from the true distribution. diffHOD-IA agreed with halotools-IA across all tests, including galaxy number counts, and the $\xi(r)$, $\omega(r)$, and $\eta(r)$ statistics for the fiducial galaxy catalog.
@@
The differentiability also allows inserting diffHOD-IA into larger differentiable inference pipelines. For example, diffHOD-IA can be integrated into simulation pipelines that incorporate a differentiable particle mesh solver such as jaxPM (Li et al., 2022), along with a differentiable halo finder like jFoF (Horowitz and Bayer, 2025). This would enable end-to-end gradient flow from cosmological, HOD, and IA parameters to the galaxy field.
The differentiable correlation function framework could also be extended to 2D-projected statistics, which are more directly comparable to observational weak lensing measurements (Van Alfen et al., 2025a). Of equal interest is the extension to higher-order statistics and field level inference methods which would impose tighter constraints on parameters. A joint inference over both HOD and IA parameters represents a natural application, using the differentiable estimators for $\xi(r)$ and $\omega(r)$. In the present work, fixing the HOD avoids degeneracies between HOD and IA parameters that arise at the 2PCF level — for instance, the satellite fraction governed by $\log M_{1}$ and $\alpha$ directly affects the relative contributions of centrals and satellites to $\omega(r)$, which could partially mimic changes in $\mu_{\rm sat}$. Such joint analyses over the full 7-dimensional parameter space would require careful treatment of these degeneracies and are reserved for future work. These developments provide a foundation for efficient, gradient-based joint inference of HOD and IA parameters in current and future weak lensing surveys.
@@ Appendix B Differentiable Galaxy Clustering Statistics
Figure 8: Validation of differentiable correlation function estimators against halotools reference implementations, averaged over 20 independent catalog realizations with error bars indicating the standard deviation across realizations. Left: Galaxy position-position correlation function $\xi(r)$. The black circles show the (non-differentiable) measurement from halotools.mock_observables.tpcf, while red squares show the differentiable estimator using HOD-derived occupation probability weights. Center: Galaxy position-orientation correlation function $\omega(r)$, comparing halotools.mock_observables.ed_3d (black) with our differentiable estimator (red). Right: Galaxy orientation-orientation correlation function $\eta(r)$, comparing halotools.mock_observables.ee_3d (black) with our differentiable estimator (red). Both $\omega(r)$ and $\eta(r)$ show excellent agreement between implementations, with $\eta(r)$ exhibiting larger statistical fluctuations due to galaxy shape noise. The differentiable estimators enable gradient-based inference: $\xi(r)$ gradients flow through galaxy occupation weights from HOD parameters, while $\omega(r)$ and $\eta(r)$ gradients flow through orientation vectors from IA parameters.